OpenStudy (anonymous):

Find the extreme values of the function and where they occur. g(x)= x^2lnx

OpenStudy (zarkon):

find g' set equal to zero solve for x

OpenStudy (anonymous):

x[2ln(x)+1]=0

OpenStudy (zarkon):

so x=0 or [2ln(x)+1]=0

OpenStudy (zarkon):

can you solve this?

OpenStudy (anonymous):

I need a little help

OpenStudy (zarkon):

subtract 1 then divide by 2

OpenStudy (zarkon):

[2ln(x)+1]=0 2ln(x)=-1 ln(x)=-1/2

OpenStudy (zarkon):

so x is ...

OpenStudy (anonymous):

thats actually the part i needed help with. I had ln(x)= -1/2

OpenStudy (zarkon):

\[\ln(x)=-1/2\] \[e^{\ln(x)}=e^{-1/2}\] \[x=e^{-1/2}\]

OpenStudy (anonymous):

So is that an extreme value?

OpenStudy (zarkon):

that is a critical number

OpenStudy (anonymous):

Alright, so how do I find the extreme values? Sorry I'm not very good at this.

OpenStudy (zarkon):

plug the critical number back into the original function to get the value...to test if it is a min/max/ or neither you need to use the first or second derivative test.

OpenStudy (anonymous):

Ok I got -.183939

OpenStudy (zarkon):

ok...now you need to see if it is a max or min...

OpenStudy (anonymous):

How do I do the first derivative test?

OpenStudy (zarkon):

plug numbers that are close to you critical number (above and below) into the derivative to see if the function is increasing or decreasing.

OpenStudy (anonymous):

I tried plugging in -.5 and 0 but you cant take the ln of 0 or a negative

OpenStudy (zarkon):

try .5 and 1 you are picking numbers about the critical number e^(-1/2) which is approx .60653

OpenStudy (anonymous):

so .5 is negative and 1 is positive

OpenStudy (zarkon):

therefore you have a min

OpenStudy (anonymous):

a min at e^-1/2?

OpenStudy (zarkon):

yes the min is located at x=e^{-1/2}

OpenStudy (anonymous):

great! thank you